# Error Propagation Mechanisms and Compensation Strategies for Quantized Diffusion Models

### v2

@@ 1 Introduction
DMs (Podell et al., 2023; Chen et al., 2024) have established themselves as highly effective deep generative frameworks across diverse domains, including image synthesis (Ho et al., 2020), video creation (Yang et al., 2024) and image translation (Sasaki et al., 2021) etc. Compared with conventional SOTA generative adversarial networks (GANs), DMs exhibit superior stability, free from the common pitfalls of model collapse and posterior collapse, which ensures more reliable and diverse output generation. Although diffusion models demonstrate remarkable capabilities in generating high-fidelity and diverse images, their substantial computational and memory overhead hinders widespread adoption. This complexity primarily stems from two factors: first, the reliance on complex deep neural networks (DNNs) for noise estimation; second, the requirement for an iterative progressive denoising process to maintain synthesis quality, which may involve up to 1,000 iterative steps, substantially increasing computational demands.
@@
To address the substantial computational demands of noise estimation in diffusion models, researchers employ quantization techniques to accelerate inference across all denoising steps. Depending on whether they require fine-tuning, quantization methods can be categorized into Quantization-Aware Training (QAT) (Wu et al., 2020) and PTQ (Cai et al., 2020). QAT necessitates retraining neural networks with simulated quantization and hyperparameter optimization, which introduces significant computational overhead and deployment complexity, making it unsuitable for compute-intensive diffusion model training. In contrast, PTQ directly derives quantization correction coefficients through calibration data statistics, thus avoiding the high-cost retraining process inherent to diffusion models. Although PTQ has been widely studied in traditional vision tasks such as image classification and object detection (Bhalgat et al., 2020), it faces many different challenges in diffusion models. The architectural characteristics and training strategies of diffusion models inherently lead to the widespread presence of outliers in weight distributions, while activation values exhibit step-wise distributional variations across time steps (Li et al., 2023a). These properties pose significant challenges for quantization by inducing substantial step-wise error. Furthermore, the iterative denoising mechanism amplifies error accumulation across successive steps, where quantization errors progressively accumulate during the sampling trajectory, ultimately degrading generation fidelity. Recent advancements in PTQ (Wu et al., 2024; Li et al., 2023a; Zhao et al., 2024b) have predominantly focused on minimizing quantization errors at individual denoising steps. However, these approaches systematically neglect the critical analysis of error propagation dynamics throughout the iterative sampling trajectory. Consequently, current solutions (Li et al., 2024b) remain constrained to 4-bit quantization of both weights and activations while maintaining acceptable quality degradation, as cumulative errors across sequential denoising stages fundamentally limit lower-bit quantization viability.
@@
Contrarily, our work focuses on the quantization error propagation problem in diffusion models and proposes the first timestep-aware cumulative error compensation scheme, called TCEC. First, we construct an error propagation equation by taking the DDIM (Song et al., 2020) as a paradigmatic case, presenting the field’s inaugural closed-form solution for cumulative error. However, direct computation of cumulative errors proves computationally infeasible. Subsequently, through the implementation of timestep-sensitive online rapid estimation for per-step quantization error, we achieve a notable simplification of the computational complexity inherent in cumulative error modeling. Finally, as shown in Figure 1, we incorporate a cumulative error correction term in each generation step, dynamically mitigating errors induced by quantization. In summary, our contributions are:
@@
Notably, TCEC maintains orthogonality to existing state-of-the-art PTQ algorithms (Li et al., 2024b; Zhao et al., 2024a) that minimize per-step quantization errors. Additionally, TCEC originates from rigorous theoretical derivation, and while its derivation process uses DDIM as an example, it is equally applicable to other solvers such as DPM++ (Lu et al., 2022).
@@ 2.1 Diffusion Model
Diffusion models are a family of probabilistic generative models that progressively destruct real data by injecting noise, then learn to reverse this process for generation, represented notably by denosing diffusion probabilistic models (DDPMs) (Ho et al., 2020). DDPM is composed of two Markov chains of $T$ steps. One is the forward process, which incrementally adds Gaussian noises into real sample $x_{0}\in q(x_{0})$, In this process, a sequence of latent variables $x_{1:T}=[x_{1},x_{2},...,x_{T}]$ are generated in order and the last one $x_{T}$ will approximately follow a standard Gaussian:
@@
where $p(\mathbf{x}_{T})\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, with $\mu_{\theta}(\mathbf{x}_{t},t)$ denotes the noise estimation model, and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ denotes the variance for sampling which can be fixed to constants (Luo, 2022). The denoising process, constrained by the Markov chain, requires a huge number of iterative time steps in DDPM. DDIM generalizes the diffusion process to non-Markovian processes, simulating the diffusion process with fewer steps. It has replaced DDPM as the mainstream inference strategy. Our work focuses on accelerating the inference of the noise estimation model in DDIM, with a training-free PTQ process.
@@ 2.2 Post-Training Quantization
The noise estimation models such as UNET (Podell et al., 2023) or Transformer (Chen et al., 2024; Yang et al., 2024; Zheng et al., 2024) exhibit high computational complexity, rendering the sampling of diffusion models computationally expensive. PTQ transforms the weights and activations of the full-precision model into a low-bit format, enabling the model’s inference process to utilize the integer matrix multiplication units on the target hardware platform and accelerating the computational process (Jacob et al., 2018). Prior studies, such as PTQD (He et al., 2023) and Q-DM (Li et al., 2023b), have explored the application of quantization techniques for diffusion models. Q-Diffusion (Li et al., 2023a) and PTQ4DM (Shang et al., 2023) first achieved 8-bit quantization in text-to-image generation tasks. Subsequent research refined these methodologies through strategies such as sensitivity analysis (Yang et al., 2023) and timestep-aware quantization (Huang et al., 2024a; Wang et al., 2023). Among these works, MixDQ (Zhao et al., 2024b) introduces metric-decoupled sensitivity analysis and develops an integer programming-based method to derive optimal mixed-precision configurations. Qua2SeDiMo (Mills et al., 2025) enables high-quality mixed-precision quantization decisions for a wide range of diffusion models, from foundational U-Nets to state-of-the-art Transformers, extending the quantization lower bounds for image generation tasks to W4A8. SVDQuant (Li et al., 2024b) enhances quantization performance by integrating fine-grained quantization with singular value decomposition (SVD)-based weight decomposition, achieving W4A4 quantization while maintaining acceptable quality degradation. Meanwhile, ViDiT-Q (Zhao et al., 2024a) further explores quantization for video generation tasks (Yang et al., 2024; Zheng et al., 2024), achieving W8A8 and W4A8 with negligible degradation in visual quality and metrics. These works minimize per-step errors through more precise quantization approximations at the layer level, yet overlook the error propagation in the diffusion process. Particularly in video generation tasks, which require a greater number of inference steps, the issue of error propagation becomes exacerbated, and accumulated errors across sequential denoising stages fundamentally constrain the feasibility of lower-bit quantization. Although a few studies (Chu et al., 2024; Yao et al., 2024) have acknowledged the error propagation issue and attempted solutions, they lack rigorous theoretical analysis. Their experimental designs rely on ad-hoc fitting to specific tasks and are only validated on simple datasets such as CIFAR10.
@@
Figure 2: Qualitative visual results comparison. Prompt1: An alien octopus floats through a protal reading a newspaper. Prompt2: A middle-aged woman of Asian descent, her dark hair streaked with silver, appears fractured and splintered, intricately embedded within a sea of broken porcelain. The porcelain glistens with splatter paint patterns in a harmonious blend of glossy and matte blues, greens, oranges, and reds, capturing her dance in a surreal juxtaposition of movement and stillness. Her skin tone, a light hue like the porcelain, adds an almost mystical quality to her form.
@@ 3.1 Error Propagation Mechanisms
$x_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}x_{t}+\left(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\right)\mu_{\theta}(x_{t},t)$ (3)
@@
$\tilde{\mu}_{\theta}(x_{t},t)=\mu_{\theta}(x_{t},t)+\varepsilon_{t}$ (4)
@@
$\tilde{x}_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\tilde{x}_{t}+\left(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\right)\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$ (5)
@@
$x_{t-1}+\delta_{t-1}=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}(x_{t}+\delta_{t})+\left(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\right)\left(\mu_{\theta}(x_{t}+\delta_{t},t)+\varepsilon_{t}\right)$ (6)
@@ 3.2 Simplify Computational Complexity
Approximation 1 For a well-trained diffusion model, it is insensitive to local changes in the input, which implies that we can ignore the Jacobian term: $J_{x_{t}}\approx 0$.
@@ 3.3 Timestep-Aware Compensation
$\frac{\partial\mathcal{L}}{\partial K_{t,i}}=-2\sum_{j=1}^{H}\sum_{k=1}^{W}\left[(1-K_{t,i})\tilde{\mu}_{t,i,j,k}-\mu_{t,i,j,k}\right]\tilde{\mu}_{t,i,j,k}+2\lambda_{1}K_{t,i}=0$ (16)
@@ 4 Experiments
Table 1: Quantization Performance Comparison of Different Models. SDXL and SDXL-Turbo generate at $512^{2}$ resolution, while PixArt achieves $1024^{2}$. Evaluation metrics include FID (distribution distance) (Parmar et al., 2024), IR (human preference) (Xu et al., 2023), LPIPS (perceptual similarity) (Zhang et al., 2018), and PSNR (numerical fidelity against 16-bit references) (Li et al., 2024b).
@@ 4.1 Implementation Details
Quantization Scheme. SVDQuant (Li et al., 2024b) introduces an additional low-rank branch that can mitigate quantization challenges in both weights and activations, establishing itself as a new benchmark for PTQ algorithms. In this study, we build our quantization strategy upon it by integrating cumulative error correction mechanisms. This approach ensures that performance comparisons remain unaffected by operator-level quantization configurations. In the 8-bit configuration, our approach employs per-token dynamic quantization for activations and per-channel weight quantization, complemented by a low-rank auxiliary branch with a rank of 16. For the 4-bit configuration, we apply per-group symmetric quantization to both activations and weights, using a low-rank branch with rank 32 and setting the group size to 64. All nonlinear activation and normalization layers are not quantized, meanwhile the inputs of linear layers in adaptive normalization are kept in 16 bits. To comprehensively evaluate the effectiveness of TCEC, we conducted comparative experiments with recent SOTA quantization algorithms across diverse generation tasks.
Image Generation Evaluation. We benchmark TCEC using SDXL (Podell et al., 2023), SDXL-Turbo (Podell et al., 2023) and PixArt models (Chen et al., 2023; 2024) including both the UNet and DiT backbones. SDXL-Turbo uses the default configuration of 4 inference steps, while SDXL employs the DDIM sampler with 50 steps. Since PixArt utilizes the DPM++ solver, we adapted TCEC to this advanced high-order solver to demonstrate its compatibility across different solvers. The detailed validation process is presented in Appendix D. To precompute the channel-wise scaling matrices $K$ and $\lambda_{1}$ mentioned in section 3.3, we randomly sampled 1,024 prompts from COCO dataset (Chen et al., 2015) as the calibration dataset. To evaluate the generalization capability of TCEC, we sample 5K prompts from the MJHQ-30K (Li et al., 2024a) and the summarized Densely Captioned Images(sDCI) (Urbanek et al., 2024) for benchmarking.
Video Generation Evaluation. We apply TCEC to OpenSORA (Zheng et al., 2024), the videos are generated with 100-steps DDIM with CFG scale of 4.0. We evaluate the quantized model on VBench (Huang et al., 2024b) to provide comprehensive results. Following prior research (Ren et al., 2024), we evaluate video quality from three distinct dimensions using eight carefully selected metrics. Aesthetic Quality and Imaging Quality focus on assessing the quality of individual frames, independent of temporal factors. Subject Consistency, Background Consistency, Motion Smoothness, and Dynamic Degree measure cross-frame temporal coherence and dynamics. Finally, Scene Consistency and Overall Consistency gauge the alignment of the video with the user-provided text prompt. We collected 128 samples from WebVid (Nan et al., 2024) as the calibration dataset to calculate $K$ and $\lambda_{1}$.

### v3

@@ 1 Introduction
DMs (podell2023sdxl; chen2024pixart) have established themselves as highly effective deep generative frameworks across diverse domains, including image synthesis (ho2020denoising), video creation (yang2024cogvideox) and image translation (sasaki2021unit) etc. Compared with conventional SOTA generative adversarial networks (GANs), DMs exhibit superior stability, free from the common pitfalls of model collapse and posterior collapse, which ensures more reliable and diverse output generation. Although diffusion models demonstrate remarkable capabilities in generating high-fidelity and diverse images, their substantial computational and memory overhead hinders widespread adoption. This complexity primarily stems from two factors: first, the reliance on complex deep neural networks (DNNs) for noise estimation; second, the requirement for an iterative progressive denoising process to maintain synthesis quality, which may involve up to 1,000 iterative steps, substantially increasing computational demands.
@@
To address the substantial computational demands of noise estimation in diffusion models, researchers employ quantization techniques to accelerate inference across all denoising steps. Depending on whether they require fine-tuning, quantization methods can be categorized into Quantization-Aware Training (QAT) (wu2020easyquant) and PTQ (cai2020zeroq). QAT necessitates retraining neural networks with simulated quantization and hyperparameter optimization, which introduces significant computational overhead and deployment complexity, making it unsuitable for compute-intensive diffusion model training. In contrast, PTQ directly derives quantization correction coefficients through calibration data statistics, thus avoiding the high-cost retraining process inherent to diffusion models. Although PTQ has been widely studied in traditional vision tasks such as image classification and object detection (bhalgat2020lsq+), it faces many different challenges in diffusion models. The architectural characteristics and training strategies of diffusion models inherently lead to the widespread presence of outliers in weight distributions, while activation values exhibit step-wise distributional variations across time steps (qdiffusion). These properties pose significant challenges for quantization by inducing substantial step-wise error. Furthermore, the iterative denoising mechanism amplifies error accumulation across successive steps, where quantization errors progressively accumulate during the sampling trajectory, ultimately degrading generation fidelity. Recent advancements in PTQ (wu2024ptq4dit; qdiffusion; mixdq) have predominantly focused on minimizing quantization errors at individual denoising steps. However, these approaches systematically neglect the critical analysis of error propagation dynamics throughout the iterative sampling trajectory. Consequently, current solutions (svdqunat) remain constrained to 4-bit quantization of both weights and activations while maintaining acceptable quality degradation, as cumulative errors across sequential denoising stages fundamentally limit lower-bit quantization viability.
@@
Contrarily, our work focuses on the quantization error propagation problem in diffusion models and proposes the first timestep-aware cumulative error compensation scheme, called TCEC. First, we construct an error propagation equation by taking the DDIM (song2020denoising) as a paradigmatic case, presenting the field’s inaugural closed-form solution for cumulative error. However, direct computation of cumulative errors proves computationally infeasible. Subsequently, through the implementation of timestep-sensitive online rapid estimation for per-step quantization error, we achieve a notable simplification of the computational complexity inherent in cumulative error modeling. Finally, as shown in Figure 1, we incorporate a cumulative error correction term in each generation step, dynamically mitigating errors induced by quantization. In summary, our contributions are:
@@
Notably, TCEC maintains orthogonality to existing state-of-the-art PTQ algorithms (svdqunat; vidit) that minimize per-step quantization errors. Additionally, TCEC originates from rigorous theoretical derivation, and while its derivation process uses DDIM as an example, it is equally applicable to other solvers such as DPM++ (lu2022dpm).
@@ 2.1 Diffusion Model
Diffusion models are a family of probabilistic generative models that progressively destruct real data by injecting noise, then learn to reverse this process for generation, represented notably by denosing diffusion probabilistic models (DDPMs) (ho2020denoising). DDPM is composed of two Markov chains of $T$ steps. One is the forward process, which incrementally adds Gaussian noises into real sample $x_{0}\in q(x_{0})$, In this process, a sequence of latent variables $x_{1:T}=[x_{1},x_{2},...,x_{T}]$ are generated in order and the last one $x_{T}$ will approximately follow a standard Gaussian:
@@
where $p(\mathbf{x}_{T})\sim\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, with $\mu_{\theta}(\mathbf{x}_{t},t)$ denotes the noise estimation model, and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ denotes the variance for sampling which can be fixed to constants (luo2022understanding). The denoising process, constrained by the Markov chain, requires a huge number of iterative time steps in DDPM. DDIM generalizes the diffusion process to non-Markovian processes, simulating the diffusion process with fewer steps. It has replaced DDPM as the mainstream inference strategy. Our work focuses on accelerating the inference of the noise estimation model in DDIM, with a training-free PTQ process.
@@ 2.2 Post-Training Quantization
The noise estimation models such as UNET (podell2023sdxl) or Transformer (chen2024pixart; yang2024cogvideox; zheng2024open) exhibit high computational complexity, rendering the sampling of diffusion models computationally expensive. PTQ transforms the weights and activations of the full-precision model into a low-bit format, enabling the model’s inference process to utilize the integer matrix multiplication units on the target hardware platform and accelerating the computational process (jacob2018quantization). Prior studies, such as PTQD (he2023ptqd) and Q-DM (li2023q), have explored the application of quantization techniques for diffusion models. Q-Diffusion (qdiffusion) and PTQ4DM (shang2023post) first achieved 8-bit quantization in text-to-image generation tasks. Subsequent research refined these methodologies through strategies such as sensitivity analysis (yang2023efficient) and timestep-aware quantization (huang2024tfmq; wang2023towards). Among these works, MixDQ (mixdq) introduces metric-decoupled sensitivity analysis and develops an integer programming-based method to derive optimal mixed-precision configurations. Qua2SeDiMo (qua2sedimo) enables high-quality mixed-precision quantization decisions for a wide range of diffusion models, from foundational U-Nets to state-of-the-art Transformers, extending the quantization lower bounds for image generation tasks to W4A8. SVDQuant (svdqunat) enhances quantization performance by integrating fine-grained quantization with singular value decomposition (SVD)-based weight decomposition, achieving W4A4 quantization while maintaining acceptable quality degradation. Meanwhile, ViDiT-Q (vidit) further explores quantization for video generation tasks (yang2024cogvideox; zheng2024open), achieving W8A8 and W4A8 with negligible degradation in visual quality and metrics. These works minimize per-step errors through more precise quantization approximations at the layer level, yet overlook the error propagation in the diffusion process. Particularly in video generation tasks, which require a greater number of inference steps, the issue of error propagation becomes exacerbated, and accumulated errors across sequential denoising stages fundamentally constrain the feasibility of lower-bit quantization. Although several studies (chu2024qncd; timestep) have recognized the issue of error propagation and attempted to propose solutions, their efforts are focused on ”single-step error source suppression” (e.g., TAC decomposes input/noise errors, while QNCD targets noise in the embedding layer). These works neither elucidate the relationship between single-step errors and cumulative errors nor validate the model performance beyond small academic datasets such as CIFAR10.
@@
Figure 2: Qualitative visual results comparison. Prompt1: An alien octopus floats through a portal reading a newspaper. Prompt2: A middle-aged woman of Asian descent, her dark hair streaked with silver, appears fractured and splintered, intricately embedded within a sea of broken porcelain. The porcelain glistens with splatter paint patterns in a harmonious blend of glossy and matte blues, greens, oranges, and reds, capturing her dance in a surreal juxtaposition of movement and stillness. Her skin tone, a light hue like the porcelain, adds an almost mystical quality to her form.
@@ 3.1 Error Propagation Mechanisms
$\begin{split}x_{t-1}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}x_{t}\\ &\quad+\left(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\right)\mu_{\theta}(x_{t},t)\end{split}$ (3)
@@
$\tilde{\mu}_{\theta}(\tilde{x}_{t},t)=\mu_{\theta}(\tilde{x}_{t},t)+\varepsilon_{t}$ (4)
@@
$\begin{split}\tilde{x}_{t-1}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\tilde{x}_{t}\\ &\quad+\left(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\right)\tilde{\mu}_{\theta}(\tilde{x}_{t},t)\end{split}$ (5)
@@
$\begin{split}x_{t-1}+\delta_{t-1}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}(x_{t}+\delta_{t})\\ &\quad+\biggl(\sqrt{1-\alpha_{t-1}}-\frac{\sqrt{\alpha_{t-1}(1-\alpha_{t})}}{\sqrt{\alpha_{t}}}\biggr)\\ &\quad\cdot\bigl(\mu_{\theta}(x_{t}+\delta_{t},t)+\varepsilon_{t}\bigr)\end{split}$ (6)
@@ 3.2 Simplify Computational Complexity
Approximation 1 For a well-trained diffusion model, it is insensitive to local changes in the input, which implies that we can ignore the Jacobian term: $J_{x_{t}}\approx 0$. See Appendix A for details.
@@ 3.3 Timestep-Aware Compensation
$\begin{split}\frac{\partial\mathcal{L}}{\partial K_{t,i}}&=-2\sum_{j=1}^{H}\sum_{k=1}^{W}\bigl[(1-K_{t,i})\tilde{\mu}_{t,i,j,k}-\mu_{t,i,j,k}\bigr]\tilde{\mu}_{t,i,j,k}\\ &\quad+2\lambda_{1}K_{t,i}=0\end{split}$ (16)
@@ 4 Experiments
Table 1: Quantization Performance Comparison of Different Models. SDXL and SDXL-Turbo generate at $512^{2}$ resolution, while PixArt achieves $1024^{2}$. Evaluation metrics include FID (distribution distance) (fid), IR (human preference) (ir), LPIPS (perceptual similarity) (lpips), and PSNR (numerical fidelity against 16-bit references) (svdqunat).
@@ 4.1 Implementation Details
Quantization Scheme. SVDQuant (svdqunat) introduces an additional low-rank branch that can mitigate quantization challenges in both weights and activations, establishing itself as a new benchmark for PTQ algorithms. In this study, we build our quantization strategy upon it by integrating cumulative error correction mechanisms. This approach ensures that performance comparisons remain unaffected by operator-level quantization configurations. In the 8-bit configuration, our approach employs per-token dynamic quantization for activations and per-channel weight quantization, complemented by a low-rank auxiliary branch with a rank of 16. For the 4-bit configuration, we apply per-group symmetric quantization to both activations and weights, using a low-rank branch with rank 32 and setting the group size to 64. All nonlinear activation and normalization layers are not quantized, meanwhile the inputs of linear layers in adaptive normalization are kept in 16 bits. To comprehensively evaluate the effectiveness of TCEC, we conducted comparative experiments with recent SOTA quantization algorithms across diverse generation tasks.
Image Generation Evaluation. We benchmark TCEC using SDXL (podell2023sdxl), SDXL-Turbo (podell2023sdxl) and PixArt models (chen2023pixart; chen2024pixart) including both the UNet and DiT backbones. SDXL-Turbo uses the default configuration of 4 inference steps, while SDXL employs the DDIM sampler with 50 steps. Since PixArt utilizes the DPM++ solver, we adapted TCEC to this advanced high-order solver to demonstrate its compatibility across different solvers. The detailed validation process is presented in Appendix D. To precompute the channel-wise scaling matrices $K$ and $\lambda_{1}$ mentioned in section 3.3, we randomly sampled 1,024 prompts from COCO dataset (chen2015microsoft) as the calibration dataset. Appendix I provides a detailed explanation of the selection and size of the calibration dataset. To evaluate the generalization capability of TCEC, we sample 5K prompts from the MJHQ-30K (li2024playground) and the summarized Densely Captioned Images(sDCI) (urbanek2024picture) for benchmarking.
Video Generation Evaluation. We apply TCEC to OpenSORA (zheng2024open), the videos are generated with 100-steps DDIM with CFG scale of 4.0. We evaluate the quantized model on VBench (huang2024vbench) to provide comprehensive results. Following prior research (ren2024consisti2v), we evaluate video quality from three distinct dimensions using eight carefully selected metrics. Aesthetic Quality and Imaging Quality focus on assessing the quality of individual frames, independent of temporal factors. Subject Consistency, Background Consistency, Motion Smoothness, and Dynamic Degree measure cross-frame temporal coherence and dynamics. Finally, Scene Consistency and Overall Consistency gauge the alignment of the video with the user-provided text prompt. We collected 128 samples from WebVid (nan2024openvid) as the calibration dataset to calculate $K$ and $\lambda_{1}$.